x squared - y squared = 39
x - y = 3

I don't know how to set up this problem to find a solution.
Thanks!

Answer by Nate(3500)

x - y = 3 ~> x = y + 3
Plug:
x^2 - y^2 = 39
(y + 3)^2 - y^2 = 39
y^2 + 6y + 9 - y^2 = 39
6y + 9 = 39
6y = 30
y = 5
Plug:
x - y = 3
x - 5 = 3
x = 8
Answers: (8,5)